Fermi National Laboratory


Accelerator Update

Wednesday February 4
The day shift began with Operations monitoring store 3217, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 8:24 AM, operators noticed that the output for a Linac quadrupole (L:QPS104) power supply was dropping to zero. Experts investigated.

Switchyard experts began a study period at 8:50 AM.

At 9:56 AM, Pbar stopped stacking. Operators investigated and discovered several racks of equipment without power. The Duty Electrician had power restored within thirty minutes. Stacking resumed at 10:24 AM.

A Main Injector (MI) expert began a study period at 10:02 AM.

At 11:44 AM, ComEd lost a 345KV high voltage line between Lockport and Lisle that caused a Fermi power glitch. An Accumulator RF station (ARF1) tripped, two cryo system valves closed, some Linac quadrupole power supplies tripped, and the Tevatron Electron Lens (TEL) tripped off. Everything reset okay except for L:QPS104. Technicians eventually got the power supply working.

Another power glitch hit the lab at 12:42 PM. Many of the same items tripped off. ComEd said a high voltage line in Elmhurst tripped this time.

QPS104 got noisy again at 1:43 PM. A Linac tech had it back online twenty minutes later.

Linac experts switched to the I- Source at 2:31 PM.

Both Switchyard and MI studies were completed by 7:50 PM.

At 8:32 PM, a Pbar expert began a reverse proton study period. This halted stacking for the rest of the evening shift. But on the bright side, with no stacking operators were able to increase the beam intensity for MiniBooNE.

Thursday February 5
The midnight shift began with Operations monitoring store 3217, with Pbar conducting studies, and with MiniBooNE taking beam.

At 2:40 AM, Pbar completed their studies and resumed stacking.

At 3:26 AM, a kicker power supply (I:KPS1C) tripped off for the first of four times in an hour. At 4:38 AM, operators called in an expert. When this kicker trips off it halts stacking.

After numerous trips of a MiniBooNE power supply (E:HV8645) operators called in an expert. He arrived and began investigating the problem at 7:18 AM.

At 8:07 AM, Operations terminated store 3217.

Operations began preparing the Tevatron (TeV) for an access at 8:14 AM for TeV maintenance.

Operations resumed sending beam to MiniBooNE at 8:45 AM.

At 9 AM, both D-zero (D0) and the Collider Detector at Fermilab (CDF) experimenters accessed their collision halls. Both accesses were completed within thirty minutes.

Techs completed their TeV maintenance at 10:41 AM.

Switchyard studies began at 11:48 AM and ran for two hours.

At 11:49 AM, Operations put the TeV through a dry squeeze.

Operations put the TeV into shot setup at 1:39 PM and at 3:54 PM established store 3219. Its initial luminosity was 58.93E30. This set a New Luminosity Record.

At 4:03 PM, techs began work on a kicker power supply (I:KPS1C). This work held off stacking until 4:40 PM.

Switchyard studies resumed at 4:53 PM.

A Linac RF station (KRF5) had a phase problem due to its cavity temperature at 5:40 PM. This happened again at 6:22 PM and 6:38 PM. Operators called in a Linac tech that repaired a LCW relay and had the system working normally by 7:47 PM.

Friday February 6
The midnight shift began with Operations monitoring store 3219, with Pbar stacking, with Switchyard conducting studies, and with MiniBooNE taking beam.

At 3:38 AM, a mysterious problem caused Booster to trip off multiple times. Operations worked on this problem for two hours and then it disappeared as mysteriously as it had appeared. (Operators hate those kinds of problems.)

Today's Plans
The plan for today is to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 2/6/2004   email Fermilab